Solve
5/2x +1/4x =5/4+x

Answer by rfer(16322)

5/2x+1/4x=5/4+x
10/4x+1/4x=5/4+4/4x
7/4x=5/4
x=(5/4)/(7/4)
x=5/4*4/7
x=20/28
x=5/7